Problem 2: The state of stress at a point is given by 80 20 40 [a] 20 60 10|MPa 40 10 20 (a) Determine the strains using YounPlease include all parts!
Answer #1

a) Strain

\sigma =\begin{bmatrix} 80 & 20 &40\\ 20&60 &10 \\ 40 &10 &20 \end{bmatrix}

But standard representation of stress matrix is :

Ơi Try Ty:

Also,

Strain in x direction : \varepsilon _{x} = \frac{\sigma _{x}}{E} - \mu\frac{\sigma _{y}}{E}-\mu\frac{\sigma _{z}}{E}

80 103 0.25.60 2520 = 60 * 10-5

Similarly , \varepsilon _{y} = \frac{\sigma _{y}}{E} - \mu\frac{\sigma _{x}}{E}-\mu\frac{\sigma _{z}}{E}

\therefore \varepsilon _{y} =\frac{60}{10^{5}} - 0.25\frac{80}{10^{5}}-0.25\frac{20}{10^{5}} = 35 * 10^{-5}

Similarly, \varepsilon _{z} = \frac{\sigma _{z}}{E} - \mu\frac{\sigma _{x}}{E}-\mu\frac{\sigma _{y}}{E}

\therefore \varepsilon _{z} =\frac{20}{10^{5}} - 0.25\frac{80}{10^{5}}-0.25\frac{60}{10^{5}} = -15 * 10^{-5}

b) Strain Energy Density

u_{x}= \frac{1}{2}\sigma_{x} *\varepsilon_{x}

u_{x}= \frac{1}{2}(80 *10^{6})*(60*10^{-5})= 24000 J/m^{3} = 24 KJ/m^{3}

Similarly , u_{y}= \frac{1}{2}\sigma_{y} *\varepsilon_{y}

u_{y}= \frac{1}{2}(60*10^{6})*(35*10^{-5})= 10500 J/m^{3} = 10.5 KJ/m^{3}

Similarly, u_{z}= \frac{1}{2}\sigma_{z} *\varepsilon_{z}

u_{z}= \frac{1}{2}(20*10^{6})*(-15*10^{-5})= -1500 J/m^{3} = -1.5 KJ/m^{3}

C) Principal Stress

RESULTS Parameter Value Unit Principal stress-1 (01) Principal stress-2 (o) Principal stress-3 () Max shear stress-1 (Tmax Ma

d)

Shear Strain in x direction ; \phi _{x}= \frac{\tau_{xy} }{G}

where , \tau_{xy} = Shear Stress\ on \ x\ plane\ in\ y \ direction

G = Modulus \ of \ Rigidity

But G = \frac{E}{2(1+\mu )} = \frac{100}{2(1+.25 )} = 40 MPa

\therefore \phi _{xz}= 40/40 = 1 = \phi _{zx}

Similarly , \therefore \phi _{yz}= 10/40 = 0.4 = \phi _{zy}

\therefore \phi _{xy}= 20/40 = 0.5 =\phi _{yx}

Now, Standard strain matrix representation

\sigma =\begin{bmatrix} \varepsilon _{_{x}} & \frac{\phi_{xy}}{2} &\frac{\phi_{xz}}{2}\\ \frac{\phi_{yx}}{2}&\varepsilon _{_{y}} &\frac{\phi_{yz}}{2} \\ \frac{\phi_{zx}}{2} &\frac{\phi_{zy}}{2} &\varepsilon _{_{z}} \end{bmatrix}

\sigma =\begin{bmatrix} \60 * 10^{-5} & 0.25 &0.5\\ 0.25&35* 10^{-5}&0.2 \\ 0.5 &0.2 &-15* 10^{-5} \end{bmatrix}